OpenStudy (anonymous):

help with 3 questions about triangles?

OpenStudy (reemii):

sure?

OpenStudy (anonymous):

please help!

OpenStudy (reemii):

you need to remember two things for triangles: 1) cos(a) = (adjacent) / (hypothenuse) (->then sin(a) = (opposite)/(hypothenuse)) 2) tan(a) = sin(a)/cos(a)  tan(a) = sin a /cos a = opposite / adjacent = 12/16 right?

OpenStudy (reemii):

i mean tan("x") = 12/16..

OpenStudy (anonymous):

yeah 3/4

OpenStudy (reemii):

sure, and the second one is quite immediate too. what do you get?

OpenStudy (reemii):

straightforward*

OpenStudy (anonymous):

i just dont know how to do these lol...

OpenStudy (reemii):

sin(x) = opposite / hypothenuse

OpenStudy (reemii):

write the fraction, and simplify

OpenStudy (anonymous):

30/34= 15/17?

OpenStudy (reemii):

why 30?

OpenStudy (reemii):

it's the opposite to the angle x, so it's 16/34

OpenStudy (anonymous):

and the last one? how do i figure that one out

OpenStudy (reemii):

now you know x, x = 63°. They say "use the cosine". The formula is cos(63) = adj / hyp right? It turns out that in this equation, only "hyp" is unknown. Therefore you just have to isolate it.

OpenStudy (anonymous):

so how do i figure it out?

OpenStudy (reemii):

53*

OpenStudy (reemii):

write the equation for the cosine: \(\cos(53) = \frac{7}{H}\). -> \(H = \dots\). use a calculator to compute cos(53°) and you're done.

OpenStudy (anonymous):

my answer is coming out to be .6

OpenStudy (reemii):

for the cos(53) ? yes. then, what is H?

OpenStudy (anonymous):

9.29

OpenStudy (reemii):

what formula did you use?

OpenStudy (reemii):

\[\cos(53°) = \frac7H\\ \iff H = \frac{7}{cos(53°)} = \frac7{0.602} = \dots\]

OpenStudy (anonymous):

9.29?

OpenStudy (reemii):

It's not that hard: if you don't have a calculator, you can use google. it'll tell you "7/0.602 = 11.6279069767 = 11.63" more or less

OpenStudy (reemii):

@ogkat it's done.